OpenStudy (anonymous):

how to determine range of a function? example (x-2)^2 + 2

OpenStudy (aum):

What is the smallest value of (x-2)^2 ?

OpenStudy (aum):

When something is squared, it is always \(\ge 0\). So \((x-2)^2 \ge 0\). Its lowest value is 0 and it occurs when x = 2.

OpenStudy (anonymous):

how about the domain of the same function

OpenStudy (aum):

So the smallest value of \((x-2)^2 + 2\) is 2.

OpenStudy (aum):

Range is [2, infinity)

OpenStudy (anonymous):

actually something with a square root would help me better with the domain

OpenStudy (aum):

Is there any restriction that we should place on x? Is there ny value of x for which \((x-2)^2 + 2\) cannot be calculated?

OpenStudy (anonymous):

so lets say instead of (x-2)^2 it's \[\sqrt{x-2}\]

OpenStudy (aum):

Since for real values, we cannot take the square root of a number < 0, we must say \(x-2 \ge 0\) or \(x \ge 2\) is the domain.

OpenStudy (anonymous):

alright that makes sense

OpenStudy (aum):

In interval notation we would write the domain as: [2, infinity)